OpenStudy (anonymous):

Two objects of mass m1 and m2 ha ve the same KE and are both moving to the right. The same constant force F is applied the left to both masses. If m1 =4m2, the ratio of the stopping distance of m1 to that of m2 is ___?

OpenStudy (sidsiddhartha):

w=F*s=(ma)s so m is inversy proportional to s m1/m2=s2/s1 s2/s1=4